# Secure and Efficient Migration of Large Enclaves in a Data Center

### v4

@@ 6. Security Analysis
ConstMig The following attacks are prevented during the migration process:
Replay attack: An adversary might intercept messages transmitted during migration and attempt to replay them to migrate the application to a compromised destination. Any attempt to replay stale data will result in an integrity check failure during restoration (§4.6) and will lead to an immediate abort of the migration process (see Invariant 4). Replaying the data during the same migration session will result in the restore thread dropping the packet after decryption, as the restoration of the replayed page is already done (see Invariant 3).
@@
Changing the Fault Address: An attempt by a malicious OS to disrupt the execution flow by changing the faulting virtual address from $VA_{A}$ to $VA_{B}$ will not result in any security leak. Based on the manipulated faulting request, the contents of the page $VA_{B}$ will be retrieved from the source machine in an encrypted form, which is guaranteed to be genuine (Invariant 1). Since the encrypted packet contains the address $VA_{B}$5), the destination will write the page contents to address $V_{B}$, which is correct. Of course, the application will again immediately fault as the page fault for the address $VA_{A}$ is yet to be serviced. However, the point to note is that there is no security violation.
@@ 7. Evaluation
We evaluate two configurations of ConstMig: ConstMig-V1, where an enclave is initialized with the maximum of required memory and ConstMig-V2, where we leverage the EDMM capabilities of SGX and initialize with a smaller amount of committed memory and add memory later as required (§2).
We compare ConstMig with the current state-of-the-art MigSGX (Nakashima and Kourai, 2021). All other prior work (Nakatsuka et al., 2022; Liang et al., 2019; Guerreiro et al., 2020; Soriente et al., 2019) is based on a similar migration scheme (see Table 4). We also evaluate TDX– TDX with no migration due to the non-availability of two machines with TDX capability. We only report the initialization time for a CVM on the destination machine – a lower bound on the total downtime, as there will be other costs, such as save, transfer, and page faults (Intel, 2025b).
Most of the workloads in our evaluation setup either do not issue a lot of free() calls and hence are unaffected by the ratio of committed to uncommitted memory (Kumar et al., 2025). Hence, we allocate most of the workloads with $\approx$ 100 MB to 200 MB of committed memory with the exception of OpenSSL. OpenSSL, within an enclave, allocated memory, performs encryption of a file, and frees the memory. It repeats the process for every new request. Hence, it is initialized with a 458 MB of committed pages. We saw no benefits in the performance for more than 458 MB of committed memory.
@@ 7.1. Experimental Setup
We use two 2-socket systems, each with two Intel Xeon Gold 6354 CPUs (72 cores), 1TB of memory, 512GB of SSD, and 1 TB of HDD storage. The system runs the Linux kernel version 5.15. A 1 Gbps network connects the systems. Both the systems support scalable SGX with an EPC of up to 64 GB. We use Intel SGX SDX v2.13 and the SGX driver (Intel, 2019b) for our SGX setup. We use CRIU (Checkpoint and Restore In Userspace) (CRIU Project, 2025), v3.16.1, as our migration manager and use LLVM for the compiler pass. Table 2 shows the workloads used to evaluate ConstMig. For TDXevaluation, we use a machine with an Intel Xeon Gold 6526Y CPU, 256GB of memory, and 2 TB of HDD storage.
7.2. Fault Tracker v.s. Access Tracker
Table 3, shows the time taken by each step in Fault and Access Trackers (resp.) for 4 KB data and 4 MB data, respectively. For 4 KB and 4 MB data, Access Tracker outperforms Fault Tracker by $35\%$ and $93\%$, respectively. The performance improvement for Access Tracker grows with an increase in the size of the data to be restored. Fault Tracker operates at a 4 KB granularity; hence, restoring 4 MB data will require $\approx$ 1,000 faults, signals, and ECALLs. Whereas, Access Tracker can potentially use the MemArr to restore the complete 4 MB region using a single ECALL. Since Access Tracker outperforms Fault Tracker, we use it for the rest of the experiments.
Table 3. Time taken to service faults of data size 4 KB and 4 MB in the restored enclave is calculated using Fault Tracker (FT) and Access Tracker (AT). All times are in  ${\mu}s$.
4 KB 4 MB
Ops FT ( ${\mu}s$) AT ( ${\mu}s$) FT ( ${\mu}s$) AT ( ${\mu}s$)
AES + SHA 33 33 33,792 3,629
ECALL 8 8,192
Signal 8 8,397
Fault 2 2,253
Total 51(1.5$\times$) 33 52,634 (14$\times$) 3,629
Figure 8. Lower is better. Total downtime for MigSGX, TDX, ConstMig-V1, and ConstMig-V2. The main cause of downtime in MigSGX is the transfer time, and in TDXand ConstMig-V1 it is the CVM/enclave init time.
7.3. Improvement in the Total Downtime
@@
The asynchronous heap migration in ConstMig, which removes it from the critical path, translates to a significant reduction in the total downtime of an enclave. ConstMig-V1 reduces the total downtime when compared with MigSGX by up to 84%. ConstMig-V2 outperforms ConstMig-V1 by up to 96%. The additional performance gains in ConstMig-V2 are due to a reduction in the time of initialization of a fresh enclave on the destination machine (low amount of committed memory). Figure 8 shows an overview of the improvement in downtime of ConstMig-V1 and ConstMig-V2 compared to MigSGX.
7.4. End-to-End Performance (Single Migration)
@@
Efficient migration also helps in improving the end-to-end performance, as the application is resumed almost instantly on the destination. In terms of the total end-to-end performance of the application being migrated (see Table 4). ConstMig-V1 outperforms MigSGX by up to $\approx$12.61%. ConstMig-V2 outperforms MigSGX by up to $\approx$12.95%. The end-to-end performance gain is due to a quick resumption of the application on the destination machine for ConstMig-V1 and ConstMig-V2. Note that in MigSGX, once data restoration is complete, the application runs without any overhead. However, in ConstMig, even though the application starts responding almost immediately, there are performance overheads due to a background data restore and network page faults. In spite of this, ConstMig outperforms in terms of end-to-end performance.
Table 4. End-to-end performance evaluation with MigSGX, ConstMig-V1 (CM-v1), and ConstMig-V2 (CM-v2). See Figure 8 for total downtime.
Setting Committed Mem. Init. Time Total Exec. time Impr. vs. MigSGX
BFS MigSGX 5 GB $\approx$8.1s 293.4s
CM-v1 5 GB $\approx$8.1s 256.4s 12.61%
CM-v2 105 MB $\approx$0.2s 255.4s 12.95%
B-Tree MigSGX 2 GB $\approx$4.3s 89.2s s
CM-v1 2 GB $\approx$4.3s 82.5s 7.51%
CM-v2 105 MB $\approx$0.2s 80.5s 9.75%
Hash- Join MigSGX 2 GB $\approx$4.3s 352.2s
CM-v1 2 GB $\approx$4.3s 320.2s 9.09%
CM-v2 105 MB $\approx$0.2s 318.2s 9.65%
Key- Value MigSGX 5.3 GB $\approx$8.3s 448.2s
CM-v1 5.3 GB $\approx$8.3s 416.1s 7.16%
CM-v2 105 MB $\approx$0.2s 413.6s 7.72%
Open- SSL MigSGX 2 GB $\approx$4.3s 122.1s
CM-v1 2 GB $\approx$4.3s 114.3s 6.39%
CM-v2 458 MB $\approx$1s 110.7s 9.34%
Page- Rank MigSGX 2 GB $\approx$4.3s 183.2s
CM-v1 2 GB $\approx$4.3s 176.4s 3.71%
CM-v2 105 MB $\approx$0.2s 169.1s 7.70%
SVM MigSGX 2 GB $\approx$4.3s 406.2s
CM-v1 2 GB $\approx$4.3s 361.7s 10.96%
CM-v2 205 MB $\approx$0.4s 358.1s 11.84%
7.4.1. Overhead of Additional Memory Checks
@@
We need to ensure low overhead when the application is not migrating (or is done migrating). In the access tracker, the additional checks return immediately. There is undoubtedly a small overhead involved; however, a lot of this gets hidden because of the predictability of the branch and the unused resources of out-of-order Intel CPUs. The total overhead of doing so is limited to a maximum of 0.16%. In the fault tracker, there is no performance overhead as there are no faults generated once the migration is done (or not yet started).
7.4.2. Performance w.r.t. No Migration
Figure 9 compares the performance overheads in MigSGX, ConstMig-V1, and ConstMig-V2 compared to baseline execution (application executing in SGX, but no migration). Compared to the baseline setting, ConstMig-V2 causes an average performance overhead of 34% in our testing setting. MigSGX causes a performance overhead of 49% over the baseline setting.
Figure 9. Figure showing end-to-end performance overhead due to migration with MigSGX, ConstMig-V1 (OM-v1), and ConstMig-V2 (OM-v2). Baseline is SGX + no migration.
7.5. Throughput Results (Key-Value Store)
@@ 8. Related Work
Table 5 presents an overview of the prior work in this area. Gu et al. (Gu et al., 2017) and Yuhala et al. (Yuhala et al., 2021) use the stop-and-copy method to migrate SGX applications from one machine to another. The former focused on the efficient transfer of keys using a remote attestation process, while the latter focused on resuming machine learning applications in case of an application crash. Nakashima and Kourai (Nakashima and Kourai, 2021) and Nakatsuka et al. (Nakatsuka et al., 2022), in their work MigSGX and CTR, respectively, point out that the enclave size can be in GBs, and using the stop-and-copy method might require a large amount of memory. However, their main contribution is in reducing the size of the unsecure buffer required to transfer the enclave data from one machine to another by employing pipelined memory transfer. Their scope is far limited as compared to our work that looks at full end-to-end migration.
@@
MigGSX (Nakashima and Kourai, 2021) Intel Stop-and-Copy & Pipeline GBs
@@
ConstMig Intel Post Copy GBs
@@ 9. Conclusion
In this paper, we present ConstMig– an efficient migration mechanism for large SGX enclaves. We show the scalability challenges of current approaches that still use the decade-old stop-and-copy migration mechanism to migrate SGX enclaves. We presented an efficient and secure saving and restoring mechanism that enables asynchronous migration of enclave data. This essentially makes the downtime of an enclave constant during migration irrespective of its size. Furthermore, we proposed two novel techniques to prevent the use-before-restore issue in a restored enclave. One technique overloads the SGX page fault handler, and the second augments each memory access with an additional check. We implemented ConstMig and evaluated it with a set of real-world applications. Our results show that ConstMig reduces total downtime by 96% for large enclaves compared to the state-of-the-art.
@@ A.4. Monitoring PageMap
Appendix B SGX’s flexible execution framework
As discussed, Intel SGX divides an application into secure and unsecure portion. The secure portion executes in an enclave inside SGX and incurs the overhead due to security guarantees of SGX. The untrusted portion however, suffers no additional performance overhead. Application executing in a CVM on Intel TDX incurs performance overhead for the complete application.
Figure 11 show the end-to-end performance of Key-Value workload for varying amount of sensitive data – for 100% everything executes in SGX and, say, for 0% only 60% data is in SGX, rest in the untrusted region outside SGX. With 100% of data as secure, SGX has a performance overhead of $\approx$12compared to TDX. However, with 20% of data as secure, SGX outperforms TDX by $\approx$ 9% and approaches the performance of No-SGX.
Appendix C Intel TDX Live migration – Init Time
Intel TDX has hardware support to perform live migration. However, to the best of our knowledge, the performance of TDX The two documents from Intel either use simulated numbers (Wang, 2022) or propose TDX live migration as an “upcoming feature” (Fuhry, 2024). Although the Git repository for the TDX module and Rust-based migration clearly states that it supports live migration (Intel, 2025c, b). Kuvaiskii et al.(Kuvaiskii et al., 2024) and Misono et al.(Misono et al., 2024) report that the initialization time of a CVM increases with an increase in the size of memory allocated – with 256 GB memory taking $\approx 24$ seconds. However, in our experiments on TDX with Canonical TDX (Canonical, 2025), we observe an initialization time of $\approx$ 26 seconds, irrespective of the CVM size.
EDMM and TDX-Shim: As per our understanding, Intel TDX does not support EDMM – a feature that allows for dynamic addition of secure data pages to an enclave (§2.2). With EDMM, the initialization time of an enclave can be significantly improved(McKeen et al., 2016). Hence, a CVM on Intel TDX must EACCEPT all the pages before it can be initialized – which will increase the boot time. However, they have an optimization built into TDX called lazy addition of the pages (Containers, 2025). Here, a background thread keep adding pages to the CVM after it has booted up. The effect of this lazy addition can be seen in the difference of memset() operation for the first-time allocation + use and the second allocation + use. If we attempt to allocate memory and use it immediately after a CVM has booted up – as will be the case for the destination CVM during a migration, we see a performance overhead of 87% (see Figure 12).
Experiment: We perform a boot, allocate-use, free, allocate-use experiment to measure the performance overhead due to lazy addition on the first allocation. We boot up a CVM with 20 GB of memory, wait exactly delay ms, and attempt to allocate and use (memset) the allocated memory. We vary delay and report the time for allocation and memset operation.
Figure 12. Performance of alloc and memset with varying the amount of delay of operation after a CVM boot. Due to lazy addition of pages, an alloc and memset immediately after a CVM boot results in a performance overhead of 100%, compared to an alloc and memset done after waiting $\approx$ 10 seconds for a 20 GB of CVM.
Appendix D Migration in Virtualized Environments
With the advent of lightweight virtualization techniques, customers deploy their stateful applications on docker containers and microVMs instead of standalone applications (Stubbs et al., 2015; Jia and Witchel, 2021). We evaluate the performance of ConstMig using three different types of sandbox environments: Docker containers (Fortune, 2023), Firecracker’s microVMs (Agache et al., 2020), and Qemu VM (qemu, 2023). Our aim is to show the improvement in migrating a secure portion of the containers. Hence, we configure our benchmarks with a small untrusted memory footprint. Using a large amount of unsecure memory will overshadow the cost of migrating SGX enclave in the presence of sandboxes with high memory footprints.
Table 6. Performance evaluation of total downtime (DT) for ConstMig on an academic cloud setup. Here, DT includes downtime due to sandbox transfer and SGX enclave transfer. Memory in MB. We perform live VM migration.
Downtime (seconds)
Benchmark Input Size (MB) Docker MicroVM VM
BFS 9 4.81 12.56 16.4
B-Tree 64 12.98 12.56 28.69
Hash-Join 50 11.01 12.56 28.05
Key-Value 80 15.08 12.59 29.26
OpenSSL 80 15.08 12.59 29.26
Pagerank 80 15.08 12.59 29.26
SVM 24 7.32 12.56 27.05
Table 6 shows the total downtime of an application during migration from one machine to another in the cloud. A Docker container and a Firecracker’s microVM officially support the stop-and-copy migration technique (doc, [n. d.]). In the case of Docker, we only transfer the anonymous memory pages over the network; therefore, the downtime is proportional to the input size. On the other hand, the downtime of a microVM is almost the same for all the benchmarks. This is because a microVM’s snapshot contains a memory file whose size is equal to the amount of memory that has been configured for the microVM. As a result, the cost of transferring the memory file over the network is the same for all benchmarks. The downtime for a VM increases by up to 3.7$\times$ compared to other sandbox environments because the resident memory size increases by 14.25$\times$. If these workloads are executed in a non-virtualized setting, the total downtime in ConstMig-V1 is $\approx$ 1s and in ConstMig-V2 is $\approx$ 0.2s, as the downtime is not impacted by the additional overheads due to the virtualization.

### v5

@@ 5. Security Analysis
Table 1 lists the key components of ConstMig along with their trust level and the data or access pattern accessible to them. The design and implementation of ConstMig ensures protection against the following attacks.
Replay attack: An adversary might intercept messages transmitted during migration and attempt to replay them to migrate the application to a compromised destination. Any attempt to replay stale data will result in an integrity check failure during restoration (§3.5.3) and will lead to an immediate abort of the migration process (see Invariant 4). Replaying the data during the same migration session will result in the restore thread dropping the packet after decryption, as the restoration of the replayed page is already done (see Invariant 3).
Entity Trusted? Access
Operating system Page fault sequence
Migration manager (CRIU) Page fault sequence
Trusted Library (ConstMigLib) Everything
Save thread (save_th) Everything
Restore thread (restore_th) Everything
Table 1. Entities involved in migration and their trust level.
@@
Changing the Fault Address: An attempt by a malicious OS to disrupt the execution flow by changing the faulting virtual address from $VA_{A}$ to $VA_{B}$ will not result in any security leak. Based on the manipulated faulting request, the contents of the page $VA_{B}$ will be retrieved from the source machine in an encrypted form, which is guaranteed to be genuine (Invariant 1). Since the encrypted packet contains the address $VA_{B}$6(a)), the destination will write the page contents to address $V_{B}$, which is correct. Of course, the application will again immediately fault as the page fault for the address $VA_{A}$ is yet to be serviced. However, the point to note is that there is no security violation.
5.1. Fault Tolerance
Fault Tolerance A migration process may fail for several reasons: source crashes, destination crashes, network failures, etc. What happens in such a situation depends on the security guarantees of the environment. To handle such cases, we extend the proposal by  (Gu et al., 2017). The original enclave (OE) and restored enclave (RE) can communicate securely via remote attestation. ❶ RE, on receiving $M_{k}$, sends a message to OE. The save_th requires the message to start saving encrypted data in BBuff. If OE does not receive the message, it times out and resumes. Even if RE is executing, it will be a fresh instance of the enclave as it will not receive any data from OE. ❷ If it does receive the message, it is aware of two things: RE has received $M_{k}$ and has resumed. OE cannot be resumed now. If, for any reason, RE crashes, the application instance is lost. ❸ Once save_th is done, it zeros all the secure pages and kills OE (since all the data has been saved).
@@ 6. Evaluation
We evaluate two configurations of ConstMig: ConstMig-V1, where the enclave is initialized with the maximum required memory upfront, and ConstMig-V2, which leverages SGX’s EDMM capabilities to initialize with a smaller committed memory footprint and dynamically add pages as needed (Dhanraj et al., 2025; Kumar et al., 2025). We compare ConstMig with the current state-of-the-art MigSGX (Nakashima and Kourai, 2021). All other prior work (Nakatsuka et al., 2022; Liang et al., 2019; Guerreiro et al., 2020; Soriente et al., 2019) is based on a similar migration scheme. We also evaluate TDX– TDX with no migration due to the non-availability of two machines with TDX capability. We only report the initialization time for a CVM on the destination machine – a lower bound on the total downtime, as there will be other costs, such as save, transfer, and page faults (Intel, 2025b).
extitWorkloads: Table 2 lists the workloads used for evaluation. Most of the workloads in our evaluation setup either do not issue a lot of free() calls and hence are unaffected by the ratio of committed to uncommitted memory (Kumar et al., 2025; Dhanraj et al., 2025). Hence, we initialize most of the workloads with $\approx$ 100 MB to 200 MB of committed memory with the exception of OpenSSL. OpenSSL, within an enclave, allocates memory, performs encryption of a file, and frees the memory. It repeats the process for every new request. Hence, it is initialized with $\approx 460$ MB of committed memory.
@@
(a) Total initialization time
(b) Total downtime
(c) Total execution time – single migration
Figure 10. Figure showing end-to-end performance overhead due to migration with MigSGX, TDX, ConstMig-V1 (CM-v1), and ConstMig-V2 (CM-v2). Baseline is SGX + no migration.
@@ 6.1. Experimental Setup
We use two 2-socket systems, each with two Intel Xeon Gold 6354 CPUs (72 cores), 1TB of memory, 512GB of SSD, and 1 TB of HDD storage. The system runs the Linux kernel version 5.15. A 1 Gbps network connects the systems. Both the systems support scalable SGX with an EPC of up to 64 GB. We use Intel SGX SDX v2.13 and the SGX driver (Intel, 2019b) for our SGX setup. We use CRIU (Checkpoint and Restore In Userspace) (CRIU Project, 2025) as our migration manager and use LLVM for the compiler pass. Table 2 shows the workloads used to evaluate ConstMig. For TDXevaluation, we use a machine with an Intel Xeon Gold 6526Y CPU, 256GB of main memory, and 2 TB of HDD storage. Our changes are mainly limited to the SGX driver and SDK.
6.2. Evaluation – Single Migration
Here, we evaluate the impact of different migration schemes when an application is migrated just once.
6.2.1. Total Downtime
@@
Figure 10(a) and Figure 10(b) show the total enclave and CVM initialization time and the total downtime experienced by the workloads for a single migration for MigSGX, TDX, ConstMig-V1, and ConstMig-V2. The asynchronous heap migration in ConstMig and TDX, which removes it from the critical path, translates to a significant reduction in the total downtime of an enclave/CVM. ConstMig-V1 reduces the total downtime when compared with MigSGX and ConstMig-V1 by up to 99.4% and 96.2%, respectively. The additional performance gains in ConstMig-V2 are due to a reduction in the time of initialization of a fresh enclave on the destination machine by reducing the committed memory in ConstMig-V2. The impact on the final execution time for the workloads is shown in Figure 10(c). ConstMig-V2 outperforms MigSGX and ConstMig-V1 by an average of $\approx 10\%$ and $\approx 2\%$, respectively. However, the performance difference widens for a long running application with multiple migration (see §6.3).
6.2.2. End-to-End Performance (Single Migration)
@@
Efficient migration also helps in improving the end-to-end performance, as the application is resumed almost instantly on the destination. In terms of the total end-to-end performance of the application being migrated. ConstMig-V1 outperforms MigSGX by up to $\approx$12.61%. ConstMig-V2 outperforms MigSGX by up to $\approx$12.95%. The end-to-end performance gain is due to a quick resumption of the application on the destination machine for ConstMig-V1 and ConstMig-V2. Note that in MigSGX, once data restoration is complete, the application runs without any overhead. However, in ConstMig, even though the application starts responding almost immediately, there are performance overheads due to a background data restore and network page faults. In spite of this, ConstMig outperforms in terms of end-to-end performance.
6.3. At-scale – Multiple Migrations
Here, we measure the performance impact of different migration mechanisms when a long-running B-Tree workload with a working set of 40 GB is migrated multiple times.
Figure 11. Performance overhead in B-Tree with different number of migrations.
As shown in Figure 11, the total execution performance improvement for ConstMig-V2 scales with the number of migrations. For a single migration, ConstMig-V2 outperforms MigSGX, TDX, ConstMig-V1 by $\approx 11\%$, $2.6\%$, and 2.32.3%, respectively. However, for a total of 16 migrations, ConstMig-V2 outperforms MigSGX, TDX, ConstMig-V1 by $\approx 79.6\%$, $23.1\%$, and 48.648.6%, respectively. The scaling in performance improvement is due to the low downtime of the ConstMig-V2 whereas others have a large downtime which adds up with each migration. Note that the migrations are done one after another, there are no “nested” migrations.
6.4. No Migration Performance
Figure 12. Figure showing performance overhead when there is no active migration. FT: Fault Tracker, AT: Access Tracker
@@ 6.4. No Migration Performance
Migrations in data centers, if done, are not a frequent operation. Hence, we need to ensure low overhead when the application is not migrating (or is done migrating). Figure 12 shows the overhead experienced by workloads with access tracker and fault tracker. In the access tracker, the additional checks return immediately. There is undoubtedly a small overhead involved; however, a lot of this gets hidden because of the predictability of the branch and the unused resources of out-of-order Intel CPUs. The performance overhead with access tracker is limited to 0.1% to 0.6%. However, with fault tracker, there is no performance overhead as there are no faults generated once the migration is done (or not yet started).
6.5. Deep Dive – Impact on Throughput
@@ 7. Related Work
Table 3 presents an overview of the prior work in this area. Gu et al. (Gu et al., 2017) and Yuhala et al. (Yuhala et al., 2021) use the stop-and-copy method to migrate SGX applications from one machine to another. The former focused on the efficient transfer of keys using a remote attestation process, while the latter focused on resuming machine learning applications in case of an application crash. Nakashima and Kourai (Nakashima and Kourai, 2021) and Nakatsuka et al. (Nakatsuka et al., 2022), in their work MigSGX and CTR, respectively, point out that the enclave size can be in GBs, and using the stop-and-copy method might require a large amount of memory. However, their main contribution is in reducing the size of the unsecure buffer required to transfer the enclave data from one machine to another by employing pipelined memory transfer. Their scope is far more limited as compared to our work that looks at full end-to-end migration.
@@
MigSGX (Nakashima and Kourai, 2021) Intel Stop-and-Copy & Pipeline GBs
@@
ConstMig Intel Post-copy GBs
@@ 8. Conclusion
In this paper, we present ConstMig– an efficient migration mechanism for large SGX enclaves. We first highlight the scalability limitations of existing approaches, which still rely on the decade-old stop-and-copy migration technique. To address these challenges, we design a secure and efficient save-and-restore mechanism that enables asynchronous migration of enclave data, ensuring constant downtime during migration, regardless of enclave size.
Furthermore, we introduce two novel techniques to prevent the use-before-restore issue in restored enclaves: (i) overloading the SGX page-fault handler and (ii) augmenting each memory access with an additional integrity check. We implement ConstMig and evaluate it using real-world applications. Our results show that ConstMig reduces total downtime by up to 96% for large enclaves compared to the state-of-the-art.